OpenStudy (anonymous):

Please help me.. c: What is one of the solutions to the following system of equations? y2 + x2 = 53 y − x = 5 (−10, −5) (−7, −2) (7, 2) (5, 10)

HanAkoSolo (jamierox4ev3r):

first you can make the numbers the same y-x=5 can equal y^2-x^2=5^2 then you solve. Does this make sense? @raebaby420 ?

OpenStudy (anonymous):

Make numbers the same?

OpenStudy (akashdeepdeb):

@Jamierox4ev3r How is yix=5 y^2-x^2 = 25 ???

HanAkoSolo (jamierox4ev3r):

well thats the way i learned it... because 5^2=25 does it not?

OpenStudy (akashdeepdeb):

@raebaby420 All you've got to do is use some identities! :D do you know the value of (x+y)^2 and (x-y)^2 ?

OpenStudy (anonymous):

Whatttt?

OpenStudy (anonymous):

Math is a different language to me and all you guys are doing is confusing me even more :c

OpenStudy (akashdeepdeb):

Haha..Sorry! Know this =! Inthis question we can eliminate the bottom 2 choices!

OpenStudy (akashdeepdeb):

C) (x,y) = (7,2) but y-x = 5 And 2-7 = -5 [SO ELIMINATE THIS] D) cannot be eliminated yet!

OpenStudy (akashdeepdeb):

Do you know any Identity?

OpenStudy (anonymous):

No..

OpenStudy (anonymous):

Math is not my subject.. science im good at... math is me flunking!

OpenStudy (akashdeepdeb):

Try this? ~ http://www.math.com/tables/algebra/polynomials.htm

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

s whats the awnser